ANALYSIS OF HIGH-RESOLUTION INFRARED SPECTRA OF ^{11}BF_{3}

TONY MASIELLO, \underline{THOMAS A. BLAKE}, Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RLO 1830); ARTHUR MAKI, 15012 24th Ave. S.E., Mill Creek WA, 98012.

We have been engaged in the measurement and analysis of high-resolution infrared spectra of enriched samples of ^{11}BF_{3}. The Bruker IFS 120HR Fourier transform spectrometer at the Pacific Northwest National Laboratory (PNNL) facilities has been used to obtain measurements that range in resolution from 0.0015 to 0.0035 cm^{−1} with pathlengths of up to 32 m. The fundamental infrared active transitions and many hot bands, overtones and combination bands have been investigated. A multiple reflection White cell was used to identify a number of weak states not directly visible from the ground state. One of these states, 3ν_{4}, visible as the hot band 3ν_{4} −2ν_{4}, is involved in a strong Fermi resonance with the ν_{3} fundamental and was directly measured for the first time. A rotational perturbation was found to couple the 2ν_{2} and ν_{1} + ν_{4} vibrations through a ΔK = 2, Δl = −1 interaction constant. The infrared inactive ν_{1} symmetric-stretching mode was characterized by two independent routes, one using the transitions (110^{0}0^{0}) − (000^{0}0^{0}), A′′_{2} − A′_{1} and (110^{0}0^{0}) − (100^{0}0^{0}), A′′_{2} − A′_{1}, and the other using the transitions (001^{1}0^{0}) −(000^{0}0^{0}), E′ − A′_{1} and (001^{1}0^{0}) − (100^{0}0^{0}), E′ − A′_{1}. Since the rovibrational parameters for the ν_{1} state are not equilibrium parameters (not parameters from the bottom of the potential well), they are slightly different for ^{11}BF_{3} and ^{10}BF_{3}. Rovibrational constants will be presented for a number of combination/overtone states as well as the Hamiltonian used to derive them.